# No-hair and almost-no-hair results for static axisymmetric black holes and ultracompact objects in astrophysical environments

### v1

@@ Acknowledgements.
GGM thanks Béatrice Bonga, Alejandro Jiménez Cano, Ariadna Ribes Metidieri, Miguel Sánchez, Stefano Liberati, José M.M. Senovilla and Raül Vera for very enlightening discussions and Robert Bryant for helpful correspondence. Financial support was provided by the Spanish Government through the Grants No. PID2020-118159GB-C43, PID2020-118159GB-C44, PID2023-149018NB-C43 and PID2023-149018NB-C44 (funded by MCIN/AEI/10.13039/501100011033), and by the Junta de Andalucía through the project FQM219. GGM is funded by the Spanish Government fellowship FPU20/01684. CB and GGM acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033. RCR acknowledges financial support through a research grant (29405) from VILLUM fonden.
@@ Appendix A Topology of event horizons
Every cross section $S$ of the horizons considered here is a compact, orientable (since it is a boundary) surface with a positive definite metric and an axial Killing vector field $\bm{m}$. According to the Poincaré-Hopf theorem Milnor and Weaver (1997) the sum of the indices of the zeros of $\bm{m}$ should be equal to the Euler number of the cross section:
@@
with $g(S)$ being the genus of the surface $S$. The vector field $\bm{m}$ can have at most isolated zeros. Given that $\bm{m}$ is the Killing vector generating axial symmetry, we have that the index of those zeros is always $+1$, as they need to behave locally around the zero as rotations since the flow of $\bm{m}$ leaves the zero as a fixed point (Fig. 13)
Figure 13: The index of the axial Killing vector field $\bm{m}$ is always $+1$ since it generates always orbits that can be mapped to a circle $S^{1}$, generating a rigid rotation around the zero of the vector, which is fixed by the action of the isometry group.
@@
One may wonder under which conditions a toroidal cross-section is possible, specifically whether there exists a type of matter content that can distort the black hole horizon to such an extent that its topology is that of a torus rather than a sphere. First, we have to take into account the fact that we need violations of the energy conditions in the external region according to the theorem by Hawking (see Prop. 9.3.2 of  Hawking and Ellis (2023)), or the one by Galloway in which some global assumptions are relaxed Galloway (1993).
@@
We can now consider the analytical continuation to the region inside the black hole. Actually, we can perform a change of coordinates to consider the maximal analytic extension of this metric. We first reparametrize the time coordinate by making the transformation $t/2m=2\tilde{t}$. Then, we can make the change of coordinates
$R^{2}-T^{2}=4r^{2},$ $T/R=\tanh(t/2),$ (67)
@@
Similar results could be obtained for any distorted black hole, which tantamount to replacing $U_{T},V_{T}$ with $U_{T}+\hat{U},V_{T}+\hat{V}$, and it is expected to find the same kind of causal structure presented here Geroch and Hartle (1982). We have restricted our analysis in the main text to the spherical black holes since the toroidal ones are not so well understood and probably not interesting for astrophysical contexts. Furthermore, there are not any explicit examples of known toroidal black holes with regular asymptotic regions (for some examples with singularities in the external regions see Thorne (1975); Peters (2008)).
@@ Appendix B Bach-Weyl rings
$U(\bm{x})=-\int d^{3}\bm{x^{\prime}}\frac{\rho(\bm{x^{\prime}})}{\absolutevalue{\bm{x}-\bm{x^{\prime}}}}.$ (70)
It is possible to analytically perform this integral and express it in terms of elliptic functions $K$ as
@@
Everywhere except for the ring where matter is located, this function constitutes a harmonic function, and hence it is solution to Laplace equation. We can take this as a solution to generate a solution of Eqs. (14)-(15) of GR. This is the so-called Bach-Weyl ring Bach and Weyl (2012). Plugging $U$ into the equations for $V$, it turns out that $V$ can be found analytically Bach and Weyl (2012):
@@
$-8\frac{z^{2}+(r-r_{\text{ring}})^{2}}{z^{2}+(r+r_{\text{ring}})^{2}}\frac{z^{2}+r^{2}+r_{\text{ring}}^{2}}{z^{2}+(r+r_{\text{ring}})^{2}}\left[K^{\prime}\left(\frac{4rr_{\text{ring}}}{z^{2}+(r+r_{\text{ring}})^{2}}\right)\right]^{2}\Bigg{]}$ (72)
This function $V_{\text{ring}}(r,z)$ obeys all the requirements of regularity of the metric: it approaches zero quadratically in $r$ as we approach the $z$-axis while it goes smoothly to zero at spatial infinity. This solution can be simply translated in the $z$ axis by making the shift $z\rightarrow z-z_{0}$ above.
@@
in Fig. 7 for the ring lying on the same plane as the black hole $z_{0}=0$ and a value $z_{0}\neq 0$.
@@ C.1 Curzon metric
with $M$ a real parameter. We need to distinguish the situation in which $m$ is positive and the situation in which $m$ is negative. To characterize the metric, it is convenient for our purposes the area of the constant $R=\sqrt{r^{2}+z^{2}}$ equipotential surfaces. For that purpose, it is convenient to change coordinates to spherical coordinates
@@
[…]
which explicitly reads:
[…]
which upon performing the integral over $\varphi$ and a rearranging of terms reads
$A(R)=2\pi R^{2}e^{2M/R-M^{2}/(2R^{2})}\int_{-1}^{1}d(\cos\theta)e^{M^{2}\cos^{2}\theta/(2R^{2})},$ (84)
and by parity arguments we can multiply by two and integrate between 00 and $1$:
[…]
We can absorb $M$ into the radial coordinate by performing the change $R\rightarrow MR$, since it only corresponds to a choice of units and we find:
[…]
@@
$\mathcal{K}=-\frac{8M^{2}e^{\frac{2M\left(M\sin^{2}(\theta)-2R\right)}{R^{2}}}\left(3R^{2}\left(M^{2}(\cos(2\theta)-3)+4MR-2R^{2}\right)-2M^{3}\sin^{2}(\theta)(M-3R)\right)}{R^{10}},$ (88)
@@
$\lim_{R\rightarrow 0}\mathcal{K}(R,0)=\lim_{R\rightarrow 0}\mathcal{K}(R,\pi)=0,$ (90)
It is quite surprising that as we approach the (at least coordinate) singularity located at $z=r=0$, the Kretschmann may vanish or not, depending on the direction through which we approach it. This is an archetypal example of directional singularity. We can depict the equipotential surfaces and the surfaces of constant Kretschmann scalar for the sake of comparison with their Schwarzschild counterpart.
Furthermore, we can take a look at the area function to see that the constant $R$ surfaces decrease in area as we decrease the value of the coordinate $R$ until we reach the surface at which $R\simeq 0.538905M$. From that point on, we have that the area of the surfaces increases without a bound. In the limit $R\rightarrow 0$, the area goes to infinity Stachel (1968).
Figure 16: We are depicting here the area function as a function of $R$ for $M=1$ in blue and $M=-1$ in red. Although for $M=1$ the area decreases as we decrease $R$ until we reach a minimum $R\simeq 0.538905$ and it increases as we decrease the value of $R$ from that value on. This contrasts with the area for $M=-1$ which is a monotonically decreasing function.
All the discussion until now has been about the nature of the constant $t$ slices of the metric, i.e., an analysis of the spacetime from the point of view of approaching the singularity through spacelike directions. However, from the point of view of general relativity, null and timelike geodesics are the most important ones since they are relevant from the point of view of the physical observers.
Regarding the surface gravity, we have that it can be obtained as
@@
If we plug in the functions for the Curzon metric, we find that the leading order behaviour at $r=z=0$ as we we approach it through an arbitrary direction $r=R\sin\theta$ and $z=R\cos\theta$ is given by
$\kappa^{2}(\theta)=\frac{M^{2}e^{\frac{M(M(-\cos(2\theta))+M-8R)}{2R^{2}}}}{R^{4}}+\text{subdominant terms}.$ (93)
@@
$\kappa_{\text{z-axis}}=\lim_{z\rightarrow 0}\frac{Me^{-\frac{2M}{\left|z\right|}}}{z^{2}}=0,$ (95)
as it is to be expected, given that the Kretschmann remains finite, and there is a piece in the Kretschmann scalar that contains derivatives of $\kappa$ along the directions transverse to the horizon.
@@
First of all, it is easy to see that we have that the limit
@@
Figure 17: We depict here in the left panel the constant redshift surface in the $(r,z)$ coordinate for the Curzon solution with negative mass and the Kretscmann scalar on the right panel. We normalize the graphics so that the maximum is of ($\order{1}$ ).
The singularity is such that it cannot be hit by any timelike geodesic. Furthermore, it cannot be hit by any reasonable observer, we follow here the discussion in Chakrabarti et al. (1983). An observer will be described by a trajectory that is a timelike curve $\gamma$, with $\xi^{a}$ its unit tangent vector pointing toward the future, which is not necessarily a geodesic, and hence it does not need to have a vanishing acceleration $a^{b}=\xi^{c}\nabla_{c}\xi^{b}$. A reasonable observer means that the integrated acceleration along the trajectory is not infinite, namely:
@@
$J^{c}=-\xi^{b}\nabla_{b}\left(m\xi^{b}\right),$ (98)
@@
An integration of this, demonstrates the statement about the equivalence between the existence of $m$ and the finite acceleration. The spacetime that we are considering is static, hence we have the timelike Killing vector field with components $t^{a}$, hence we have $\nabla_{(a}t_{b)}$. As a consequence of this, we have the conserved quantity along geodesics that we can identify with the effective energy:
@@
Expanding the derivative of $E$ we have $\xi^{b}\nabla_{b}E=-a_{b}t^{b}-\xi^{b}\xi^{c}\nabla_{b}t_{c}$. The last term vanishes since it is contracted with a symmetric tensor and it vanishes through the Killing equation. We have then:
@@
since in our conventions the timelike character of $t^{b}$ means that $t^{b}t_{b}\leq 0$. This leads to |$\absolutevalue{\xi^{b}\nabla_{b}E}\leq aE$. If we integrate this inequality now along the curve, we conclude that $E$ needs to be finite along the curve also. We can use the triangular inequality applied to the definition of the energy leads to
@@
Finally, the area in this case is a monotone increasing function of $R$, approaching 00 as we approach $R=0$, as it can be seen in Fig. 16.
@@ C.2 Higher-multipole configurations
We can compute the Kretschmann scalar of these functions for different multipolar configurations and all of them show a behaviour that is qualitatively similar to the Curzon metric: there exist directions for approaching the $r=z=0$ point giving rise to a curvature singularity, and others that do not.

### v3

@@ Acknowledgements.
GGM thanks Béatrice Bonga, Alejandro Jiménez Cano, Ariadna Ribes Metidieri, Miguel Sánchez, Stefano Liberati, José M.M. Senovilla and Raül Vera for very enlightening discussions and Robert Bryant for helpful correspondence. Financial support was provided by the Spanish Government through the Grants No. PID2020-118159GB-C43, PID2020-118159GB-C44, PID2023-149018NB-C43 and PID2023-149018NB-C44 (funded by MCIN/AEI/10.13039/501100011033), and by the Junta de Andalucía through the project FQM219. GGM is funded by the Spanish Government fellowship FPU20/01684. CB and GGM acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033. RCR acknowledges financial support through a research grant (29405) from VILLUM fonden. LJG also acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC).
@@ Appendix A Topology of event horizons
Every cross section $S$ of the horizons considered here is a compact, orientable (since it is a boundary) surface with a positive definite metric, and an axial Killing vector field $\bm{m}$. According to the Poincaré-Hopf theorem Milnor and Weaver (1997) the sum of the indices of the zeros of $\bm{m}$ should be equal to the Euler number of the cross section:
@@
with $g(S)$ being the genus of the surface $S$. The vector field $\bm{m}$ can have at most isolated zeros. Given that $\bm{m}$ is the Killing vector generating axial symmetry, we have that the index of those zeros is always +1+1, as they need to behave locally around the zero as rotations since the flow of $\bm{m}$ leaves the zero as a fixed point (Fig. 12)
Figure 12: The index of the axial Killing vector field $\bm{m}$ is always +1+1 since it generates always orbits that can be mapped to a circle $S^{1}$, generating a rigid rotation around the zero of the vector, which is fixed by the action of the isometry group.
@@
One may wonder under which conditions a toroidal cross-section is possible, specifically whether there exists a type of matter content that can distort the black hole horizon to such an extent that its topology is that of a torus rather than a sphere. First, we have to take into account the fact that we need violations of the energy conditions in the external region according to the theorem by Hawking (see Prop. 9.3.2 of  Hawking and Ellis (2023)), or the one by Galloway in which some global assumptions are relaxed Galloway (1993). See also Chrusciel and Wald (1994); Jacobson and Venkataramani (1995) for further developments.
@@
We can now consider the analytical continuation to the region inside the black hole. Actually, we can perform a change of coordinates to consider the maximal analytic extension of this metric. We can make the change of coordinates
$R^{2}-T^{2}=4r^{2},$ $T/R=\tanh[t/(8m)],$ (70)
@@
Similar results could be obtained for any distorted black hole, which tantamount to replacing $U_{T},V_{T}$ with $U_{T}+\hat{U},V_{T}+\hat{V}$, and it is expected to find the same kind of causal structure presented here Geroch and Hartle (1982). In the main text, we have restricted our analysis to spherical black holes, as toroidal ones are less well understood and likely less relevant in astrophysical contexts, given that they require violations of energy conditions in the black hole’s exterior region. Furthermore, there are not any explicit examples of known toroidal black holes with regular asymptotic regions (for some examples with singularities in the external regions see Thorne (1975); Peters (2008)).
@@ Appendix B Bach-Weyl rings
$|.\displaystyle U(\bm{x})=-\int d^{3}\bm{x^{\prime}}\frac{\rho(\bm{x^{\prime}})}{\absolutevalue{\bm{x}-\bm{x^{\prime}}}}$. (73)
It is possible to analytically perform this integral and express it in terms of the elliptic function $K$ Abramowitz and Stegun (1964) as
@@
Everywhere except for the ring where matter is located, this function is a harmonic function, and hence it is solution to Laplace equation. We can take this as a solution to generate a solution of Eqs. (14)-(15) of GR. This is the so-called Bach-Weyl ring Bach and Weyl (2012). Plugging $U$ into the equations for $V$, it turns out that $V$ can be found analytically Bach and Weyl (2012):
@@
$-8\frac{z^{2}+(r-r_{\text{ring}})^{2}}{z^{2}+(r+r_{\text{ring}})^{2}}\frac{z^{2}+r^{2}+r_{\text{ring}}^{2}}{z^{2}+(r+r_{\text{ring}})^{2}}\left[K^{\prime}\left(\frac{4rr_{\text{ring}}}{z^{2}+(r+r_{\text{ring}})^{2}}\right)\right]^{2}\Bigg{]}.$ (75)
This function $V_{\text{ring}}(r,z)$ obeys all the requirements of regularity of the metric: it approaches zero quadratically in $r$ as we approach the $z$-axis, while it goes smoothly to zero at spatial infinity. This solution can be simply translated in the $z$ axis by making the shift $z\rightarrow z-z_{0}$ above.
@@
in Fig. 7 for the ring lying on the same plane as the black hole $z_{0}=0$ and a value $z_{0}\neq 0$. Notice that other solutions can be found that are generated by an infinitesimally thin ring, see for instance the “elliptic metric” found in Batic et al. (2023) through the Hankel integral transform method.
@@ C.1 Curzon metric
with $M$ a real parameter. We need to distinguish the situation in which $m$ is positive and the situation in which $m$ is negative. It is convenient to change coordinates to spherical coordinates
@@
$A(R)=\int d\Omega\sqrt{g^{(2)}}=R^{2}e^{2M/R}\int_{0}^{\pi}d\theta\int^{2\pi}_{0}d\varphi\sin\theta e^{-M^{2}\sin^{2}\theta/(2R^{2})},$ (85)
which upon performing the integral over $\varphi$ and rearranging terms reads
$A(R)=2\pi R^{2}e^{2M/R-M^{2}/(2R^{2})}\int_{-1}^{1}d(\cos\theta)e^{M^{2}\cos^{2}\theta/(2R^{2})}.$ (86)
We can absorb $M$ into the radial coordinate by performing the change $R\rightarrow|M|R$, since it only corresponds to a choice of units. Also by parity arguments we can multiply by two and integrate between 0 and $1$ to obtain
$A(R)/M^{2}=4\pi R^{2}e^{2/R-1/(2R^{2})}\int_{0}^{1}due^{u^{2}/(2R^{2})}\quad M>0,$ (87)
$A(R)/M^{2}=4\pi R^{2}e^{-2/R-1/(2R^{2})}\int_{0}^{1}due^{u^{2}/(2R^{2})}\quad M<0.$ (88)
@@
$\mathcal{K}=-\frac{8M^{2}e^{\frac{2M\left(M\sin^{2}(\theta)-2R\right)}{R^{2}}}\left(3R^{2}\left(M^{2}(\cos(2\theta)-3)+4MR-2R^{2}\right)-2M^{3}\sin^{2}(\theta)(M-3R)\right)}{R^{10}}.$ (89)
@@
$\lim_{R\rightarrow 0}\mathcal{K}(R,0)=\lim_{R\rightarrow 0}\mathcal{K}(R,\pi)=0.$ (91)
It is quite surprising that as we approach the (at least coordinate) singularity located at $z=r=0$, the Kretschmann scalar may vanish or not, depending on the direction through which we approach it. This is an archetypal example of directional singularity. We have depicted, see Fig. 3, the equipotential surfaces and the surfaces of constant Kretschmann scalar for the sake of comparison with their Schwarzschild counterpart.
Furthermore, we can take a look at the area function to see that the constant $R$ surfaces decrease in area as we decrease the value of the coordinate $R$ until we reach the surface at which $R\simeq 0.538905M$, see Fig 15. From that point on, we have that the area of the surfaces increases without a bound. In the limit $R\rightarrow 0$, the area goes to infinity Stachel (1968).
Figure 15: We are depicting here the area function as a function of $R$ for $M=1$ in blue and M=1M=-1 in red. Although for $M=1$ the area decreases as we decrease $R$ until we reach a minimum $R\simeq 0.538905$ and it increases as we decrease the value of $R$ from that value on. This contrasts with the area for M=1M=-1 which is a monotonically decreasing function.
All the discussion until now has been about the nature of the constant $t$ slices of the metric, i.e., an analysis of the spacetime from the point of view of approaching the singularity through spacelike directions. However, from the point of view of GR, null and timelike geodesics are the most important ones since they are relevant from the point of view of the physical observers.
Regarding the surface gravity, we it can be obtained as
@@
If we plug in the functions for the Curzon metric, we find that the leading order behavior at $r=z=0$ as we we approach it through an arbitrary direction $r=R\sin\theta$ and $z=R\cos\theta$ is given by
[…]
@@
$\kappa_{\text{z-axis}}=\lim_{z\rightarrow 0}\frac{Me^{-{2M}/{\left|z\right|}}}{z^{2}}=0,$ (96)
as it is to be expected, given that the Kretschmann scalar remains finite, and there is a piece in the Kretschmann scalar that contains derivatives of $\kappa$ along the directions transverse to the horizon.
@@
First of all, it is easy to see that the limit
@@
[…]
The singularity is such that it cannot be hit by any timelike geodesic. Furthermore, it cannot be hit by any reasonable observer. We follow here the discussion in Chakrabarti et al. (1983). An observer will be described by a trajectory that is a timelike curve $\gamma$, with $\xi^{A}$ its unit tangent vector pointing toward the future, which is not necessarily a geodesic, and hence it does not need to have a vanishing acceleration $a^{B}=\xi^{C}\nabla_{C}\xi^{B}$. A reasonable observer means that the integrated acceleration along the trajectory is not infinite, namely:
@@
$J^{C}=-\xi^{B}\nabla_{B}\left(m\xi^{C}\right),$ (99)
@@
An integration of this, demonstrates the statement about the equivalence between the existence of $m$ and the finite acceleration. The spacetime that we are considering is static, hence we have the timelike Killing vector field with components $t^{A}$, hence we have $\nabla_{(A}t_{B)}=0$. As a consequence of this, we have the conserved quantity along geodesics that we can identify with the effective energy:
@@
Expanding the derivative of $E$ we have $\xi^{B}\nabla_{B}E=-a_{B}t^{B}-\xi^{B}\xi^{B}\nabla_{B}t_{C}$. The last term vanishes since it is contracted with a symmetric tensor and it vanishes through the Killing equation. We have then:
@@
[…]
@@
Finally, the area in this case is a monotone increasing function of $R$, approaching 0 as we approach $R=0$, as it can be seen in Fig. 15.
@@ C.2 Higher-multipole configurations
We can compute the Kretschmann scalar of these functions for different multipolar configurations and all of them show a behavior that is qualitatively similar to the Curzon metric: there exist directions for approaching the $r=z=0$ point giving rise to a curvature singularity, and others that do not.